OpenStudy (anonymous):

Suppose that a customer service center claims that the telephone calls it receives last 114 seconds on average, with a standard deviation of 9 seconds. If you took a sample of 324 telephone calls, which of the following mean times would be within the 95% confidence interval? A. 116.5 seconds B. 115.5 seconds C. 114.5 seconds D. 117.5 seconds

OpenStudy (anonymous):

The 95% confidence interval is \[\left(114-1.96\sqrt{\frac{81}{324}},~114+1.96\sqrt{\frac{81}{324}}\right)\] or \[(113.02,~114.98)\]